OpenStudy (anonymous):

Can somebody help me with this question please and EXPLAIN it .√7x(√x-7√7) A.√7x-49√x B.√7x-49x C.x√7-x√49 D.√42x

OpenStudy (anonymous):

Dude stop

OpenStudy (anonymous):

Don't just hand over the answers

OpenStudy (anonymous):

\[\sqrt{a}*\sqrt{a}=a,\sqrt{a}*a=a \sqrt{a}\]